How many O2 molecules are needed to react with 7.75 g of S? Express your answer numerically in units of molecules.
Answer #1

Sulfur is always present in diatomic molecule form and first, it gives SO2 as a product. Hence we are considering this reaction to writing a balanced equation.

S2 (s) + 2O2(g) --------> 2SO2(g)

Thus as per the balanced equation, every S2 molecule needs 2 molecules of O2. Or we can also say that each atom of S needs 1 molecule of O2.

molar mass of S2 is 64.13g/mol,

so the number of moles in 7.75 g of S2 = 7.75/64.13 = 0.1209 moles of S2,

we know each mole of a substance contains = 6.022*10^23 molecules.

the number of molecules in 0.1209 moles of S2 = 0.1209*6.022*10^23 molecules of S2 = 0.7281*10^23 molecules of S2

Thus, The number of S atoms in 7.75 g = 2*0.7281*10^23 = 1.456*10^23 atoms of S

Now each of this atom requires 1 molecule of O2, then total molecules of O2 required are = 1.456*10^23 molecules of O2.
Answer #2
2.18 *10^23